# A note on subgroups whose quotients are quasi-lines

### v1

@@ Abstract
We study finitely generated pairs of groups $H\leq G$ such that the Schreier graph of $H$ is quasi-isometric to a line. Under this hypothesis, we show that $H$ is a virtual fiber subgroup if and only if $G$ contains infinitely many double cosets of $H$. Along the way, we prove that if a group acts essentially on a finite dimensional CAT(0) cube complex with no facing triples then it virtually surjects onto the integers with kernel commensurable to a hyperplane stabiliser.
@@ 1 Introduction
A well-known result of Hopf [Hopf] states that any finitely generated two-ended group is virtually cyclic. Analogous results exist for pairs of finitely generated groups $H\leq G$ which are two-ended in appropriate ways, subject to $H$ being “sufficiently normal”. The first such result is a theorem of Houghton stating that, if $H$ has two relative ends and infinite index in its normaliser, then there exists a finite index subgroup $G^{\prime}\leq G$ such that $H$ is a normal subgroup of $G^{\prime}$ and $G^{\prime}/H$ is infinite cyclic [Houghton, Theorem 3.7].
The number of relative ends of a subgroup $H\leq G$ was introduced by Houghton in the same paper. It can be defined as follows. Fix a finite generating set $S$ for $G$ and let $\operatorname{Cay}(G,S)$ be the Cayley graph of $G$ with respect to $S$. Given a subspace $K\subseteq\operatorname{Cay}(G,S)$, we say that a connected component of $\operatorname{Cay}(G,S)-K$ is deep if it contains points arbitrarily far from $K$. Given $R\geq 0$, let $\mathcal{N}_{R}(H)$ denote the $R$-neighbourhood of $H$. The number of relative ends of $H$ is the maximum over $R\geq 0$ of the number of $H$-orbits of deep components of $\operatorname{Cay}(G,S)-\mathcal{N}_{R}(H)$. In contrast, the number of filtered ends of $H$ is the maximum over $R\geq 0$ of the number of deep components of $\operatorname{Cay}(G,S)-\mathcal{N}_{R}(H)$. This invariant was introduced by Kropholler and Roller in [Kropholler-Roller]; a more geometric perspective can be found in [Geoghegan, Chapter 14].
If $H$ has two filtered ends it might not have two relative ends but there is a subgroup of $H$ of index at most 2 which does. This stronger assumption allows for the condition that $H$ has infinite index in its normaliser to be weakened as follows. Two subgroups $H,L\leq G$ are commensurable if $H\cap K$ has finite index in both $H$ and $K$. The commensurator of $H$ is the subgroup
@@
Clearly $N(H)\leq\operatorname{Comm}_{G}(H)$, where $N(H)$ denotes the normaliser of $H$. Kropholler and Roller showed that, if $H$ has two filtered ends and infinite index in its commensurator, then there are finite index subgroup $G^{\prime}\leq G$ and $H^{\prime}\leq H$ such that $H^{\prime}$ is a normal subgroup of $G^{\prime}$ and $G^{\prime}/H^{\prime}$ is infinite cyclic [Kropholler-Roller, Theorem 1.3].
We strengthen the condition that $H$ has two relative ends in a different way, assuming instead that the quotient $H\operatorname{\setminus}\operatorname{Cay}(G,S)$ is quasi-isometric to a line. This allows us to weaken the condition on the normaliser to the assumption that there are infinitely many distinct double cosets of $H$.
@@
Let $G$ be a finitely generated group. Suppose that $H\leq G$ is a finitely generated subgroup such that $H\operatorname{\setminus}G$ is quasi-isometric to a line. Then the following are equivalent:
@@
The descending implications of the theorem all hold without the assumption that $H\operatorname{\setminus}G$ is quasi-isometric to a line. The implications (1) $\Rightarrow$ (2) and (3) $\Rightarrow$ (4) are straightforward while (2) $\Rightarrow$ (3) is due to Yang [Yang]. In contrast, none of the ascending implications hold without some extra condition on $H$ or on $G$ (see Examples 1.3, 1.4, 1.5 below). Moreover, none of the properties (1)-(4) follow directly from the fact that $H\operatorname{\setminus}G$ is quasi-isometric to a line. This is illustrated in Section 5 using a class of groups introduced by Houghton [Houghton78].
@@
If $G$ is a hyperbolic surface group and $H\leq G$ is cyclic then $H$ is separable by [Scott78] but $H$ is not a virtual fiber subgroup. Note that $H$ also has two relative ends and two filtered ends.
@@
Burger and Mozes constructed in [Burger-Mozes] a family of finitely presented simple groups which split as an amalgamated free product $F\ast_{E}F$, where $F$ is a free group of finite rank and $E$ is a finite index subgroup of both factors. Let $G=F\ast_{E}F$ be such a group. Then, since $G$ acts on a locally finite tree with $E$ an edge stabiliser, [Hruska-Wise, Theorem 3.2] implies that $E$ has bounded packing in $G$. However, the fact that $G$ is simple prevents it from having any proper finite index subgroups, so $E$ cannot be separable.
@@
Theorem 1.1 is proved in Section 3. We in fact prove a slightly more general statement (Theorem 3.1) where we allow $H\operatorname{\setminus}G$ to be quasi-isometric to a finitely ended tree with at least two ends.
Sardar showed that, if $G$ is a nilpotent-by-polycyclic group, then every finitely generated subgroup of $G$ has bounded packing [Sardar, Theorem 3.7]. We therefore have the following corollary of Theorem 1.1.
@@
Let $G$ be a finitely generated nilpotent-by-polycyclic group. If $H\leq G$ is finitely generated then $H\operatorname{\setminus}G$ is quasi-isometric to a line if and only if $H$ is a virtual fiber subgroup.
Subgroups with quotient quasi-isometric to a line.
Part of the motivation for studying subgroups whose quotients are quasi-lines comes from previous work of the author with Damien Gaboriau [AG23, Theorem 3.10] where we showed that, if $H\leq G$ is finitely generated and $H\operatorname{\setminus}G$ is quasi-isometric to a line, then there are at most countably many intermediate subgroups $H\leq K\leq G$. In fact the result is more precise: the subgroup $H$ vanishes by the third Cantor-Bendixson derivative of the space of subgroups of $G$. Even in the case where $H$ is known to be a virtual fiber, it is the quasi-isometry to $\mathbb{Z}$ which reveals this property of $H$.
@@
The bounded packing property was introduced by Hruska and Wise in [Hruska-Wise]. Examples of subgroups which satisfy this property include, on the one hand, subgroups which are normal (this was shown by Hruska and Wise) or commensurated [Connor-Mihalik-2013], and, on the other, quasiconvex subgroups of hyperbolic groups [Sageev97] (see also [Chepoi-Dragan-Vaxes, Theorem 2]) and many of their generalisations (e.g. relatively quasiconvex subgroups of suitable relatively hyperbolic groups [Hruska-Wise], strongly quasiconvex subgroups of any finitely generated group [Tran], hierarchically quasiconvex subgroups of hierarchically hyperbolic groups [HHP]). This list is highly non-exhaustive.
@@
The main content of Theorem 1.1 is the implication $(3)\Rightarrow(1)$ and this is proved using an action of $G$ on a finite dimensional CAT(0) cube complex. This cube complex is obtained using Sageev’s construction [Sageev95] applied to the codimension one subgroup $H$ and the fact that it is finite dimensional follows from the fact that $H$ has bounded packing [Hruska-Wise]. This will allow us to use techniques developed by Caprace and Sageev in [Caprace-Sageev], some of which we recall in the next section. We will also need the following proposition, which we prove in Section 4.
@@
Let $\Gamma$ be a group and suppose that $\Gamma$ acts essentially on a finite dimensional CAT(0) cube complex $Y$ which has no facing triples. Then for any hyperplane $\operatorname{\mathfrak{h}}$ in $Y$, there are finite index subgroups $K\leq\operatorname{Stab}_{\Gamma}(\operatorname{\mathfrak{h}})$ and $\Lambda\leq\Gamma$ such that $K\unlhd\Lambda$ and $\Lambda\operatorname{\setminus}K=\mathbb{Z}$.
@@
Acknowledgements
I am very grateful to Mark Hagen and Indira Chatterji, for many helpful comments and suggestions and for carefully reading several drafts of this paper. This work was funded by a University of Bristol PhD scholarship.
@@ 2 Cube complexes and facing triples
We assume some familiarity with actions on CAT(0) cube complexes; introductory material can be found in [Chatterji-Niblo, Hruska-Wise-2014, Nica]. Recall that in [Sageev95] Sageev gave a construction which, given a pair of groups $H\leq G$ such that $H$ has at least two relative ends, produces a CAT(0) cube complex on which $G$ acts with a single orbit of hyperplanes such that the stabiliser of some hyperplane is commensurable with $H$. As was pointed out by Haglund and Wise in [Haglund-Wise, Corollary 3.1], it was observed in [Sageev97] that if $H$ is finitely generated and has bounded packing then this construction yields a finite dimensional cube complex (see also [Hruska-Wise-2014, Corollary 3.31]).
@@
We will work with the piecewise $\ell^{1}$ metric on CAT(0) cube complexes (see [Miesch]) which we denote by $d$. Let $g$ be a cubical automorphism of a CAT(0) cube complex $C$ and recall that $g$ is an isometry with respect to $d$ and the CAT(0) metric (i.e. the piecewise $\ell^{2}$ metric). A combinatorial geodesic in $C$ is a geodesic path $\ell:[0,L]\rightarrow C^{(1)}$ where $L\in\mathbb{N}$ and $\ell(i)\in C^{(0)}$ for each $i\in\mathbb{Z}\cap[0,L]$. Haglund showed in [Haglund, Theorem 6.3] that $g$ either fixes a point in $C$ or admits a combinatorial geodesic axis in the cubical subdivision $C^{\prime}$ of $C$ (see [Haglund, Definition 2.3, Lemma 4.2]). In the later case, $g$ is called hyperbolic and admits a (not necessarily unique or combinatorial) geodesic axis in $C$ (i.e. a bi-infinite geodesic on which $\operatorname{\langle}g\operatorname{\rangle}$ acts by translation).
@@
Let us recall some notions and results from [Caprace-Sageev]. Let $C$ be a finite dimensional cube complex and $G$ be a group acting by cubical automorphisms on $C$. The action is said to be essential if one (equivalently every) $G$-orbit contains points arbitrarily far from any halfspace of $G$. If $\operatorname{\mathfrak{h}}$ is a hyperplane of $C$ then we denote by $\operatorname{\mathfrak{h}}^{+},\operatorname{\mathfrak{h}}^{-}$ the halfspaces bounded by $\operatorname{\mathfrak{h}}$. For any $g\in G$ we write $g\operatorname{\mathfrak{h}}^{+}$ (respectively $g\operatorname{\mathfrak{h}}^{-}$) to denote the image of $\operatorname{\mathfrak{h}}^{+}$ (respectively $\operatorname{\mathfrak{h}}^{-}$) under $g$ and $(g\operatorname{\mathfrak{h}})^{+}$ to denote an independent choice of halfspace bounded by $g\operatorname{\mathfrak{h}}$. An element $g\in G$ is said to skewer $\operatorname{\mathfrak{h}}$ if there is some $n\in\mathbb{N}$ such that either $g^{n}\operatorname{\mathfrak{h}}^{+}\subsetneq\operatorname{\mathfrak{h}}^{+}$ or $g^{n}\operatorname{\mathfrak{h}}^{-}\subsetneq\operatorname{\mathfrak{h}}^{-}$. Caprace and Sageev show in [Caprace-Sageev, Lemma 2.3] that a element $g\in G$ skewers a hyperplane $\operatorname{\mathfrak{h}}$ if and only if $g$ is hyperbolic and one (equivalently any) axis of $g$ intersects $\operatorname{\mathfrak{h}}$ at a single point. This statement refers to geodesic axes of $g$ with respect to the CAT(0) metric but the same proof shows that the statement holds for $\ell^{1}$ geodesic axes of $g$, including combinatorial axes when they exist. The following characterisation of essentiality will be useful.
Proposition 2.1 (Caprace-Sageev, [Caprace-Sageev, Lemma 2.4, Proposition 3.2]).
@@
Lemma 2.2 (Caprace-Sageev, [Caprace-Sageev, Lemma 4.4]).

### v2

# On subgroups with narrow Schreier graphs

@@ Abstract
We study finitely generated pairs of groups $H\leq G$ such that the Schreier graph of $H$ has at least two ends and is narrow. Examples of narrow Schreier graphs include those that are quasi-isometric to finitely ended trees or have linear growth. Under this hypothesis, we show that $H$ is a virtual fiber subgroup if and only if $G$ contains infinitely many double cosets of $H$. Along the way, we prove that if a group acts essentially on a finite dimensional CAT(0) cube complex with no facing triples then it virtually surjects onto the integers with kernel commensurable to a hyperplane stabiliser.
@@ 1 Introduction
A well-known result of Hopf [Hop44] states that any finitely generated two-ended group is virtually cyclic. Analogous results exist for pairs of finitely generated groups $H\leq G$ which are two-ended in appropriate ways, subject to $H$ being “sufficiently normal”. The first such result is a theorem of Houghton stating that, if $H$ has two relative ends and infinite index in its normaliser, then there exists a finite index subgroup $G^{\prime}\leq G$ such that $H$ is a normal subgroup of $G^{\prime}$ and $G^{\prime}/H$ is infinite cyclic [Hou74, Theorem 3.7].
The number of relative ends of a subgroup $H\leq G$ was introduced by Houghton in the same paper. It can be defined as follows. Fix a finite generating set $S$ for $G$ and let $\operatorname{Cay}(G,S)$ be the Cayley graph of $G$ with respect to $S$. Given a subspace $K\subseteq\operatorname{Cay}(G,S)$, we say that a connected component of $\operatorname{Cay}(G,S)-K$ is deep if it contains points arbitrarily far from $K$. Given $R\geq 0$, let $\mathcal{N}_{R}(H)$ denote the $R$-neighbourhood of $H$. The number of relative ends of $H$ is the maximum over $R\geq 0$ of the number of $H$-orbits of deep components of $\operatorname{Cay}(G,S)-\mathcal{N}_{R}(H)$. In contrast, the number of filtered ends of $H$ is the maximum over $R\geq 0$ of the number of deep components of $\operatorname{Cay}(G,S)-\mathcal{N}_{R}(H)$. This invariant was introduced by Kropholler and Roller in [KR89]; a more geometric perspective can be found in [Geo08, Chapter 14].
If $H$ has two filtered ends it might not have two relative ends but there is a subgroup of $H$ of index at most 2 which does. This stronger assumption allows for the condition that $H$ has infinite index in its normaliser to be weakened as follows. Two subgroups $H,K\leq G$ are commensurable if $H\cap K$ has finite index in both $H$ and $K$. The commensurator of $H$ is the subgroup
@@
Clearly $N(H)\leq\operatorname{Comm}_{G}(H)$, where $N(H)$ denotes the normaliser of $H$. Kropholler and Roller showed that, if $H$ has two filtered ends and infinite index in its commensurator, then there are finite index subgroups $G^{\prime}\leq G$ and $H^{\prime}\leq H$ such that $H^{\prime}$ is a normal subgroup of $G^{\prime}$ and $G^{\prime}/H^{\prime}$ is infinite cyclic [KR89, Theorem 1.3].
We alter the condition that $H$ has two relative ends in a different way. We say that an unbounded connected graph $Y$ is narrow if for every $\mu\geq 1$ there exists an upper bound on the number of pairwise disjoint, unbounded, $\mu$-coarsely connected subspaces one can find in $Y$ (see Definition 3.1). Examples of narrow graphs include graphs with linear growth and infinite uniformly locally finite quasi-trees with finitely many ends (see Proposition 3.2). Under the assumption that the quotient $H\operatorname{\setminus}\operatorname{Cay}(G,S)$ has at least two ends and is narrow for some (equivalently any) finite $G$-generating set $S$, we can weaken the condition on the normaliser to the assumption that there are infinitely many distinct double cosets of $H$ (see Theorem 1.1 below).
@@
Let $G$ be a finitely generated group and suppose that $H\leq G$ is a finitely generated subgroup such that $H\operatorname{\setminus}G$ is narrow and has at least two ends. Then the following statements are equivalent:
@@
The descending implications of the theorem all hold without any assumptions on $H\operatorname{\setminus}G$. The implications (1) $\Rightarrow$ (2) and (3) $\Rightarrow$ (4) are straightforward while (2) $\Rightarrow$ (3) is due to Yang [Yan11]. In contrast, none of the ascending implications hold without some extra condition on $H$ or on $G$ (see Examples 1.3, 1.4, 1.5 below). Moreover, none of the properties (1)-(4) follow directly from the fact that $H\operatorname{\setminus}G$ is narrow with at least two ends, or even from the stronger assumption that $H\operatorname{\setminus}G$ is quasi-isometric to a line. This is illustrated in Section 6 using a class of groups introduced by Houghton [Hou78]. Finally, the necessity of the condition that $H$ is finitely generated is illustrated by Example 1.6 below.
@@
If $G$ is a hyperbolic surface group and $H\leq G$ is cyclic then $H$ is separable by [Sco78] but $H$ is not a virtual fiber subgroup. Note that $H$ also has two relative ends and two filtered ends.
@@
Burger and Mozes constructed in [BM00] a family of finitely presented simple groups which split as an amalgamated free product $F\ast_{E}F$, where $F$ is a free group of finite rank and $E$ is a finite index subgroup of both factors. Let $G=F\ast_{E}F$ be such a group. Then, since $G$ acts on a locally finite tree with $E$ an edge stabiliser, [HW09, Theorem 3.2] implies that $E$ has bounded packing in $G$. However, the fact that $G$ is simple prevents it from having any proper finite index subgroups, so $E$ cannot be separable.
@@
Example 1.6.
Let $G$ be the Grigorchuk group – viewed as a group of automorphisms of the 2-regular rooted tree $T$ – and recall that $G$ is a finitely generated torsion group so in particular cannot contain any virtual fiber subgroups (see e.g. [Gri05]). Given a ray $\xi$ based at the root of $T$, let $G_{\xi}$ denote the stabiliser of $\xi$. The stabiliser of each vertex of $\xi$ has finite index in $G$ so $G_{\xi}$ is an intersection of finite index subgroups and is therefore separable. For many choices of $\xi$, the Schreier graph of $G_{\xi}$ is quasi-isometric to a line (see e.g. [Gri05, page 201]) so in particular it is narrow and two-ended yet $G_{\xi}$ does not satisfy the conclusion of Theorem 1.1.
The following corollary is an immediate consequence of Theorem 1.1 and Proposition 3.2.
@@
Let $H\leq G$ be a finitely generated pair of groups such that $H\operatorname{\setminus}G$ has at least two ends and either $H\operatorname{\setminus}G$ has linear growth or $H\operatorname{\setminus}G$ is quasi-isometric to a finitely ended tree. Then $H$ is a virtual fiber subgroup if and only if $|H\operatorname{\setminus}G/H|=\infty$.
Corollary 1.8.
Let $H\leq G$ be a finitely generated pair of groups and suppose that $H\operatorname{\setminus}G$ has $3\leq e<\infty$ ends and is narrow. Then $H$ has finitely many double cosets in $G$.
Proof.
Suppose that $H$ is a virtual fiber subgroup and let $H^{\prime}\leq H$ be such that the normaliser $N(H^{\prime})$ has finite index in $G$ and $N(H^{\prime})/H^{\prime}=\mathbb{Z}$. Then $H^{\prime}\operatorname{\setminus}\operatorname{Cay}(G,S)$ is a quasi-line and there is a covering map $H^{\prime}\operatorname{\setminus}\operatorname{Cay}(G,S)\rightarrow H\operatorname{\setminus}\operatorname{Cay}(G,S)$. If $H\operatorname{\setminus}\operatorname{Cay}(G,S)$ has $\geq n$ ends then there are $n$ geodesic rays $\overline{\delta_{1}},\dots,\overline{\delta_{n}}\subseteq H\operatorname{\setminus}\operatorname{Cay}(G,S)$ with infinite Hausdorff distance from each other. If $\delta_{1},\dots,\delta_{n}\subseteq H^{\prime}\operatorname{\setminus}\operatorname{Cay}(G,S)$ are lifts of $\overline{\delta_{1}},\dots,\overline{\delta_{n}}$ respectively then for all $i\neq j$, $\delta_{i}$ and $\delta_{j}$ have infinite Hausdorff distance from each other. Since $H^{\prime}\operatorname{\setminus}\operatorname{Cay}(G,S)$ is a quasi-line, this implies that $n\in\{1,2\}$. Thus the corollary follows from Theorem 1.1. ∎
Several classes of groups $G$ have been shown to have the property that if $H\leq G$ is a subgroup such that $|H\operatorname{\setminus}G/H|<\infty$ then $|G/H|<\infty$. These include finitely generated nilpotent-by-polycyclic groups [dC06, Proposition 3.20] and branch groups [Fra21, Theorem A]. We therefore have the following corollary of Theorem 1.1.
Corollary 1.9.
Let $G$ be a finitely generated group and suppose that $G$ is either nilpotent-by-polycyclic or a branch group. If $H\leq G$ is finitely generated and $H\operatorname{\setminus}G$ is narrow with at least two ends then $H$ is a virtual fiber subgroup.
Note that nilpotent-by-polycyclic groups also have the property that every finitely generated subgroup has bounded packing [Sar15, Theorem 3.7] but they are not in general subgroup separable.
Consequences for the space of subgroups.
Let $\operatorname{Sub}(G)$ denote the set of all subgroups of $G$ equipped with the topology induced from the Cantor set $\{0,1\}^{G}$ (viewed as the power set of $G$). In previous work with Damien Gaboriau, we showed that if $G$ is finitely generated and $H\leq G$ is such that $H\operatorname{\setminus}G$ is quasi-isometric to a line then there are countably many intermediate subgroups $H\leq K\leq G$ [AG23, Proposition 15]. In the case where $H$ is finitely generated this implies that $H$ is not in the perfect kernel of $\operatorname{Sub}(G)$. More precisely, we showed that such a subgroup $H$ vanishes by the third Cantor-Bendixson derivative of $\operatorname{Sub}(G)$ [AG23, Theorem 3.10] (i.e. $\operatorname{rk}_{CB}^{\times}(G;H)\leq 3$ in the language of that paper). Combining this with the results of this paper we obtain the following.
Corollary 1.10.
Let $G$ be finitely generated and $H\leq G$ be a subgroup such that $H\operatorname{\setminus}G$ is narrow. Then there are countably many intermediate subgroups $H\leq K\leq G$. If in addition $H$ is finitely generated and $H\operatorname{\setminus}G$ has at least two ends then $\operatorname{rk}_{CB}^{\times}(G;H)\leq 3$.
Proof.
Let $H\leq K\leq G$ be an intermediate subgroup. By Proposition 3.5 either $[G:K]<\infty$ or $[K:H]<\infty$. For each $k\in\mathbb{N}$, there are finitely many subgroups of $G$ of index $k$ and there are at most countably many subgroups of $G$ containing $H$ with index $k$ so this implies that there are countably many such subgroups $H\leq K\leq G$. Suppose that $H$ is finitely generated and that $H\operatorname{\setminus}G$ has at least two ends. Theorem 1.1 implies that either $H$ is a virtual fiber subgroup or $|H\operatorname{\setminus}G/H|<\infty$. In the former case $H\operatorname{\setminus}G$ is a quasi-line so $\operatorname{rk}_{CB}^{\times}(G;H)\leq 3$ by [AG23, Theorem 3.10] and in the latter $H$ is isolated in $\operatorname{Sub}(G)$ so $\operatorname{rk}_{CB}^{\times}(G;H)=1$. ∎
@@
The bounded packing property was introduced by Hruska and Wise in [HW09]. Examples of subgroups which satisfy this property include, on the one hand, subgroups which are normal (this was shown by Hruska and Wise) or commensurated [CM13], and, on the other, quasiconvex subgroups of hyperbolic groups [Sag97] (see also [CDV17, Theorem 2]) and many of their generalisations (e.g. relatively quasiconvex subgroups of suitable relatively hyperbolic groups [HW09], strongly quasiconvex subgroups of any finitely generated group [Tra19], hierarchically quasiconvex subgroups of hierarchically hyperbolic groups [HHP23]). This list is highly non-exhaustive.
@@
The main content of Theorem 1.1 is the implication $(3)\Rightarrow(1)$ and this is proved using an action of $G$ on a finite dimensional CAT(0) cube complex. This cube complex is obtained using Sageev’s construction [Sag95] applied to the codimension one subgroup $H$ and the fact that it is finite dimensional follows from the fact that $H$ has bounded packing [HW09]. This will allow us to use techniques developed by Caprace and Sageev in [CS11], some of which we recall in the next section. We will conclude using the following proposition, which may be of independent interest.
@@
Let $\Gamma$ be a group and suppose that $\Gamma$ acts essentially on a finite dimensional CAT(0) cube complex $Y$ which has no facing triples. Then for any hyperplane $\operatorname{\mathfrak{h}}$ in $Y$, there are finite index subgroups $K\leq\operatorname{Stab}_{\Gamma}(\operatorname{\mathfrak{h}})$ and $\Lambda\leq\Gamma$ such that $K\unlhd\Lambda$ and $\Lambda/K=\mathbb{Z}$.
@@
Outline of the paper.
We recall some useful definitions and facts relating to CAT(0) cube complexes in Section 2. In Section 3, we define the notion of a narrow graph and show that graphs with linear growth and graphs that are infinite, uniformly locally finite, finitely ended quasi-trees are narrow. Theorem 1.1 is proved in Section 4, modulo Proposition 1.11 which is proved in Section 5. In Section 6, we present examples for each $n\geq 2$ of finitely generated pairs of groups $H\leq G$ such that $H\operatorname{\setminus}G$ is an $n$-ended quasi-tree with linear growth (in particular narrow) but $H$ is not a virtual fiber subgroup.
Acknowledgements.
I am very grateful to Mark Hagen and Indira Chatterji for many helpful comments and suggestions and for carefully reading several drafts of this paper. I thank Yves Cornulier and Paul-Henry Leemann for pointing out Example 1.6 and Adrien Le Boudec for asking me whether a previous version of Theorem 1.1 could be extended to include subgroups whose Schreier graphs have linear growth. This work was funded by a University of Bristol PhD scholarship.
@@ 2 Cube complexes and facing triples
We assume some familiarity with actions on CAT(0) cube complexes; introductory material can be found in [CN05, HW14, Nic04]. Recall that in [Sag95] Sageev gave a construction which, given a pair of groups $H\leq G$ such that $H$ has at least two relative ends, produces a CAT(0) cube complex on which $G$ acts with a single orbit of hyperplanes such that $H$ has finite index in the stabiliser of some hyperplane. As was pointed out by Haglund and Wise in [HW08, Corollary 3.1], it was observed in [Sag97] that if $H$ is finitely generated and has bounded packing then this construction yields a finite dimensional cube complex (see also [HW14, Corollary 3.31]).
@@
We will work with the piecewise $\ell^{1}$ metric on CAT(0) cube complexes (see [Mie14]) which we denote by $d$. Let $g$ be a cubical automorphism of a CAT(0) cube complex $C$ and recall that $g$ is an isometry with respect to $d$ and the CAT(0) metric (i.e. the piecewise $\ell^{2}$ metric). A combinatorial geodesic in $C$ is a geodesic path $\ell:[0,L]\rightarrow C^{(1)}$ where $L\in\mathbb{N}$ and $\ell(i)\in C^{(0)}$ for each $i\in\mathbb{Z}\cap[0,L]$. Haglund showed in [Hag23, Theorem 6.3] that $g$ either fixes a point in $C$ or admits a combinatorial geodesic axis in the cubical subdivision $C^{\prime}$ of $C$ (see [Hag23, Definition 2.3, Lemma 4.2]). In the later case, $g$ is called hyperbolic and admits a (not necessarily unique or combinatorial) geodesic axis in $C$ (i.e. a bi-infinite geodesic on which $\operatorname{\langle}g\operatorname{\rangle}$ acts by translation).
@@
Let us recall some notions and results from [CS11]. Let $C$ be a finite dimensional cube complex and $G$ be a group acting by cubical automorphisms on $C$. The action is said to be essential if one (equivalently every) $G$-orbit contains points arbitrarily far from any halfspace of $G$. If $\operatorname{\mathfrak{h}}$ is a hyperplane of $C$ then we denote by $\operatorname{\mathfrak{h}}^{+},\operatorname{\mathfrak{h}}^{-}$ the halfspaces bounded by $\operatorname{\mathfrak{h}}$. For any $g\in G$ we write $g\operatorname{\mathfrak{h}}^{+}$ (respectively $g\operatorname{\mathfrak{h}}^{-}$) to denote the image of $\operatorname{\mathfrak{h}}^{+}$ (respectively $\operatorname{\mathfrak{h}}^{-}$) under $g$ and $(g\operatorname{\mathfrak{h}})^{+}$ to denote an independent choice of halfspace bounded by $g\operatorname{\mathfrak{h}}$. An element $g\in G$ is said to skewer $\operatorname{\mathfrak{h}}$ if there is some $n\in\mathbb{N}$ such that either $g^{n}\operatorname{\mathfrak{h}}^{+}\subsetneq\operatorname{\mathfrak{h}}^{+}$ or $g^{n}\operatorname{\mathfrak{h}}^{-}\subsetneq\operatorname{\mathfrak{h}}^{-}$. Caprace and Sageev show in [CS11, Lemma 2.3] that a element $g\in G$ skewers a hyperplane $\operatorname{\mathfrak{h}}$ if and only if $g$ is hyperbolic and one (equivalently any) axis of $g$ intersects $\operatorname{\mathfrak{h}}$ at a single point. This statement refers to geodesic axes of $g$ with respect to the CAT(0) metric but the same proof shows that the statement holds for $\ell^{1}$ geodesic axes of $g$, including combinatorial axes when they exist. The following characterisation of essentiality will be useful.
Proposition 2.1 (Caprace-Sageev, [CS11, Lemma 2.4, Proposition 3.2]).
@@
Lemma 2.2 (Caprace-Sageev, [CS11, Lemma 4.4]).